The welding department supplies parts to the final assembly line. Management decides to implement a kanban system and has collected the following data.
  • The daily demand is 2000 units
  • The production lead time is 4 days (this includes processing time, transport time, and waiting time)
  • Management has decide to have a 1 day safety stock
  •  One container fits 400 units
 
How many kanban containers will be needed to support this system?
Answer #1

Compute the number of kanban containers in the system as below,

Number of \(\mathrm{K}\) anban \(=\frac{\mathrm{L}_{\mathrm{d}}+\mathrm{SS}}{\mathrm{n}}\)

Here.

Lead time demand, \(L_{\mathrm{d}}=4\) days \(\times 2000\) units per day \(=8,000\) units Container size, \(\mathrm{n}=400\) units Safety stock, \(\mathrm{SS}=1\) day \(\times 2,000\) units per day \(=2,000\) units

$$ \begin{aligned} \text { Number of Kanban } &=\frac{\mathrm{L}_{\mathrm{d}}+\mathrm{SS}}{\mathrm{n}} \\ &=\frac{8000+2000}{400} \\ &=25 \text { containers } \end{aligned} $$